# Toward a generalization of Lehmer’s problem to adelic curves

### v1

@@ 3. On Lehmer’s problem on adelic curves
Next, we show that the classical Lehmer’s problem generalizes naturally to the adelic curve $\mathbb{Q}(T)$.
4. Lehmer’s problem on $\mathbb{Q}(T_{1},\ldots,T_{d})$
Let $S$ be the adelic curve on $\mathbb{Q}(T)$ recalled in Paragraph 2.2. Let
$[\varphi:\psi]\in\mathbb{P}^{1}\left(\overline{\mathbb{Q}(T)}\right).$
@@ 4. Lehmer’s problem on $\mathbb{Q}(T_{1},\ldots,T_{d})$
Using [7, Lemma 3.5.4], we obtain
@@
Theorem 4.1.
(A Kronecker’s theorem)
$\{\alpha\in(\overline{\mathbb{Q}(T_{1},\ldots,T_{d})})^{\ast}\mid h_{S}(\alpha)=0\}=\{\text{roots of unity in}\ \overline{\mathbb{Q}}\}.$
@@
For simplicity, we prove the theorem when $d=1$. Let $\varphi$ and $\psi$ be two coprime polynomials in $\mathbb{Z}[T]$. We have
@@
$[1:\varphi]=[\varphi_{1}:\varphi_{2}],$
@@
If $\alpha\in\overline{\mathbb{Q}}$, then
@@
where $h$ is the standard Weil’s height of $\alpha$.
@@
Lehmer’s problem is equivalent to Lehmer’s problem on $\mathbb{Q}(T_{1},\ldots,T_{d})$.

### v2

@@ 4.3. On Lehmer’s problem on adelic curves
5. Lehmer’s problem on $\overline{\mathbb{Q}(T)}$
In this section, we show that the classical Lehmer’s problem generalizes naturally to the adelic curve $\mathbb{Q}(T)$.
Let $S$ be the adelic curve on $\mathbb{Q}(T)$ recalled in Paragraph 4.2. Let
$\varphi:\psi]\in\mathbb{P}^{1}\left(\overline{\mathbb{Q}(T)}\right).$
@@ 5. Lehmer’s problem on $\overline{\mathbb{Q}(T)}$
Using [12, Lemma 3.5.4], we obtain
@@
Theorem 5.1 (A Kronecker’s theorem).
$\{\alpha\in(\overline{\mathbb{Q}(T)})^{\ast}\mid h_{S}(\alpha)=0\}=\{\text{roots of unity in}\ \overline{\mathbb{Q}}\}.$
@@
Let $\varphi$ and $\psi$ be two coprime polynomials in $\mathbb{Z}[T]$. We have
@@
$1:\varphi]=[\varphi_{1}:\varphi_{2}],$
@@
Let $\alpha\in\overline{\mathbb{Q}}$. It is clear that
@@
where $h$ is the standard Weil height of $\alpha$.
@@
Lehmer’s problem is equivalent to Lehmer’s problem on $\mathbb{Q}(T)$.
Acknowledgement: I would like to thank the referee for their insightful comments and suggestions.